y'''' + 3y''' + 3y' + y = x*e^(-x) + x*cosx - 7 + x^(2) *e^(-x) *sinx


1
Expert's answer
2022-01-03T16:29:15-0500

The corresponding homogeneous equation


"y''' + 3y'' + 3y' + y =0"

Characteristic (auxiliary) equation


"r^3+3r^2+3r+1=0"

"(r+1)^3=0"

"r_1=r_2=r_3=-1"

The general solution of the homogeneous differential equation is


"y_h=c_1e^{-x}+c_2xe^{-x}+c_3x^2e^{-x}"

Find the partial solution of the nonhomogeneous differential equation


"y''' + 3y'' + 3y' + y =xe^{-x} - 7""y_1(x)=x^3(Ax+B)e^{-x}+C"

"y_1'(x)=(-Ax^4-Bx^3+4Ax^3+3Bx^2)e^{-x}"

"y_1''(x)=(Ax^4+Bx^3-4Ax^3-3Bx^2)e^{-x}"

"+(-4Ax^3-3Bx^2+12Ax^2+6Bx)e^{-x}"

"y_1'''(x)=(-Ax^4-Bx^3+8Ax^3)e^{-x}"

"+(6Bx^2-12Ax^2-6Bx)e^{-x}"




"+(4Ax^3+3Bx^2-24Ax^2)e^{-x}"

"+(-12Bx+24Ax+6B)e^{-x}"

Substitute


"(-Ax^4-Bx^3+8Ax^3)e^{-x}"

"+(6Bx^2-12Ax^2-6Bx)e^{-x}"


"+(4Ax^3+3Bx^2-24Ax^2)e^{-x}"

"+(-12Bx+24Ax+6B)e^{-x}"




"+3(Ax^4+Bx^3-4Ax^3-3Bx^2)e^{-x}"

"+3(-4Ax^3-3Bx^2+12Ax^2+6Bx)e^{-x}"

"+3(-Ax^4-Bx^3+4Ax^3+3Bx^2)e^{-x}"

"+(Ax^4+Bx^3)e^{-x}+C =xe^{-x} - 7"

"x^4e^{-x}:0=0"

"x^3e^{-x}:0=0"

"x^2e^{-x}:0=0"

"x^1e^{-x}:24A=1"

"x^0e^{-x}:B=0"

"x^0:C=-7"

The partial solution of the nonhomogeneous differential equation


"y''' + 3y'' + 3y' + y =xe^{-x} - 7"

is

"y_1(x)=\dfrac{x^4e^{-x}}{24}"

Find the partial solution of the nonhomogeneous differential equation


"y''' + 3y'' + 3y' + y =x\cos x"

"y_2(x)=Ax\cos x+Bx\sin x+C\cos x+D\sin x"


"y_2'(x)=A\cos x-Ax\sin x+B\sin x+Bx\cos x"

"-C\sin x+D\cos x"

"y_2''(x)=-2A\sin x-Ax\cos x+2B\cos x"

"-Bx\sin x-C\cos x-D\sin x"

"y_2'''(x)=-3A\cos x+Ax\sin x-3B\sin x"

"-Bx\cos x+C\sin x-D\cos x"

Substitute


"-3A\cos x+Ax\sin x-3B\sin x"

"-Bx\cos x+C\sin x-D\cos x"

"-6A\sin x-3Ax\cos x+6B\cos x"

"-3Bx\sin x-3C\cos x-3D\sin x"

"+3A\cos x-3Ax\sin x+3B\sin x+3Bx\cos x"

"-3C\sin x+3D\cos x"

"+Ax\cos x+Bx\sin x+C\cos x+D\sin x"

"=x\cos x"

"x\cos x: 2B-2A=1"

"x\sin x: 2B+2A=0"

"\cos x: 2D+6B-2C=0"

"\sin x: -2C-6A-2D=0"

"A=-\dfrac{1}{4}, B=\dfrac{1}{4}, C=\dfrac{3}{4}, D=0"

The partial solution of the nonhomogeneous differential equation


"y''' + 3y'' + 3y' + y =x\cos x"

is


"y_2(x)=-\dfrac{1}{4}x\cos x+\dfrac{1}{4}x\sin x+\dfrac{3}{4}\cos x"

Find the partial solution of the nonhomogeneous differential equation


"y''' + 3y'' + 3y' + y =x^2 e^{-x}\sin x"

"y_3(x)=(Ax^2+Bx+C)e^{-x}(D\cos x+E\sin x)"


"y_3'(x)=(2Ax+B)e^{-x}(D\cos x+E\sin x)"

"-(Ax^2+Bx+C)e^{-x}(D\cos x+E\sin x)"

"+(Ax^2+Bx+C)e^{-x}(-D\sin x+E\cos x)"

"y_3''(x)=2Ae^{-x}(D\cos x+E\sin x)"

"-(4Ax+2B)e^{-x}(D\cos x+E\sin x)"

"+(4Ax+2B)e^{-x}(-D\sin x+E\cos x)"

"+2(Ax^2+Bx+C)e^{-x}(D\sin x-E\cos x)"

"y_3'''(x)=-6Ae^{-x}(D\cos x+E\sin x)"

"+6Ae^{-x}(-D\sin x+E\cos x)"

"+(12Ax+6B)e^{-x}(D\sin x-E\cos x)"

"-2(Ax^2+Bx+C)e^{-x}(D\sin x-E\cos x)"

"+2(Ax^2+Bx+C)e^{-x}(D\cos x+E\sin x)"

After substitution we have


"y_3(x)=(x^2\cos x-6x\sin x-12\cos x)e^{-x}"

The general solution of the nonhomogeneous differential equation is


"y=c_1e^{-x}+c_2xe^{-x}+c_3x^2e^{-x}""+\dfrac{x^4e^{-x}}{24}-\dfrac{1}{4}x\cos x+\dfrac{1}{4}x\sin x+\dfrac{3}{4}\cos x"

"+x^2e^{-x}\cos x-6xe^{-x}\sin x-12e^{-x}\cos x"